OpenStudy (professorowl):

Evaluate. 3⋅(1−2/10^−3)+12 −3,012 −2,988 2,600 3,066

OpenStudy (mhchen):

Any ideas on how to solve it?

OpenStudy (professorowl):

I've been coming up with the same answer, but my answer is not one of the answers for the question.

563blackghost (563blackghost):

Hmmm are you sure that is how the equation is set up? Cause I get an answer that is not one of the options.

OpenStudy (professorowl):

Yeah, it is

OpenStudy (professorowl):

IS your answer 27/5?

563blackghost (563blackghost):

No I did not get that.

563blackghost (563blackghost):

First we would simplify the fraction....so we would do `1-2` and `10-3`. \(\huge{\frac{1-2}{10-3} \rightarrow \color{purple}{-\frac{1}{7}}}\) In which we would then have to multiply with `3`.... \(\huge{-\frac{1\times 3}{7} \rightarrow \color{green}{-\frac{3}{7}}}\) In which we would add `12` into the equation.... \(\huge{- \frac{3}{7} + 12 \rightarrow 11\frac{4}{7} = \color{red}{ \frac{81}{7}}}\)

OpenStudy (professorowl):

@Nnesha

563blackghost (563blackghost):

But that is not one of the options.

563blackghost (563blackghost):

I'm sorry :(

OpenStudy (professorowl):

OH!

OpenStudy (professorowl):

I typed something wrong @563blackghost

OpenStudy (professorowl):

Ok I fixed it.

563blackghost (563blackghost):

Ahhh well then xD

OpenStudy (professorowl):

it was 10^-3 not 10-3

OpenStudy (mathmale):

The problem here, 3⋅(1−2/10^−3)+12, is ambiguous; different people seem to be interpreting it in different ways. All that is clear to me is that y ou're multiplying the quantity within parentheses by 3 and then adding 12 to the product. Could you use additional parentheses to clarify what is to be done within the parentheses and in what order?

OpenStudy (mathmagicain):

3⋅(1−2/10^−3)+12=-5985

OpenStudy (mathmagicain):

PEMDAS

OpenStudy (mathmale):

Other ways in which you could present this problem include drawing it (use the Draw utility) or typing it in Equation Editor (below).

OpenStudy (mathmale):

It's to your advantage to spend just a little extra time to make your post crystal clear; it'll save you time, as well as save the time of others who would like to help you.

OpenStudy (professorowl):

 Man I suck at drawing on pc.